In 1895, the first Putting Green Championship was held. The winner’s prize money was $170. In 2016, the winner’s check was $1,370,000.

  

a.

What was the percentage increase per year in the winner’s check over this period? (Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.)

b. If the winner’s prize increases at the same rate, what will it be in 2038? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.)
Answer #1

We use the formula:
A=P(1+r/100)^n
where
A=future value
P=present value
r=rate of interest
n=time period.

a.

1,370,000=170*(1+r/100)^121

(1,370,000/170)^(1/121)=(1+r/100)

(1+r/100)=1.0772

r=1.0772-1

=7.72%(Approx).

b.A=1,370,000*(1.0772)^22

=$7029942.54(Approx).